OpenStudy (olsolneo123):

5 times 5/7 (fraction)

OpenStudy (unklerhaukus):

\[a\times\frac bc=\frac{a\times b}c\]

OpenStudy (olsolneo123):

so it would be \[\left(\begin{matrix}25 \\ 7\end{matrix}\right)\]

OpenStudy (jojo4eva):

the answer is 25/7(fraction) which makes it a mixed number 3 (4/7) as a mixed number

OpenStudy (jojo4eva):

correct @olsolneo123

OpenStudy (olsolneo123):

Thank you!:)

OpenStudy (jojo4eva):

No problems let me kno if you need help with anything else

OpenStudy (olsolneo123):

ok!

OpenStudy (olsolneo123):